# The Future of Artificial Intelligence and the Mathematical and Physical Sciences (AI+MPS)

### v2

@@ 3.2.1 AI in the Context of CHE
Chemistry and AI have long been working together, creating an exchange that is reshaping both fields. On the one hand, AI is helping chemists predict reactions, design molecules, and analyze complex datasets with speed and precision—accelerating discovery across chemical synthesis, drug development, materials science, and catalysis. On the other hand, chemistry is pushing the frontiers of AI. Theoretical chemistry has inspired some of today’s most powerful generative models, including diffusion models which were directly influenced by non-equilibrium thermodynamics. In fact, chemistry has provided some of AI’s most celebrated successes; notably, AlphaFold earned the 2024 Nobel Prize in Chemistry [^29] https://www.nobelprize.org/prizes/chemistry/2024/press-release/ for its breakthrough in protein structure prediction and protein design. These examples highlight a growing, bidirectional relationship: AI is transforming how we do chemistry, and chemistry continues to shape and challenge the future of AI itself.
@@ 3.2.2 How AI is Advancing CHE
  • Computational chemistry increasingly uses AI tools to predict molecular behavior and accelerate simulations, particularly for complex systems such as protein dynamics or excited-state reactions in a complex medium. These often involve developing surrogate models to replace expensive simulation methods based on quantum mechanics. Beyond speed-ups, AI is also aiding in automated design of coarse-grained models, identifying low-dimensional collective variables that capture essential system dynamics, and enhancing sampling of rare events that are otherwise inaccessible to conventional simulations. These different sub-disciplines of computational chemistry are now rapidly being transformed with the thoughtful use of AI, often deeply integrated with statistical mechanics.
@@
  • Interatomic potentials and atomistic simulation studies have been transformed by machine learning, enabling first-principles accuracy with far greater computational efficiency. While traditional first-principles approaches like density functional theory (DFT) are limited to small systems (hundreds of atoms) and short timescales (picoseconds), machine-learned interatomic potentials (MLIPs) make it possible to simulate tens of thousands of atoms over nanoseconds or longer. These advances provide efficient and accurate ways to predict reaction pathways and transition states, rapidly screen molecular libraries, and design novel hypothetical materials with targeted properties for experimental synthesis.
@@ 3.2.4 Opportunities in AI+CHE
  • Foundation models: Foundation models have shown promise in handling a range of chemistry tasks, from property prediction to retrosynthesis and even literature mining. Fine-tuning them for synthesis, materials, or protein design could streamline workflows, but questions remain about how well they generalize to out-of-distribution chemistry or rare reactions. The longer-term opportunity is to build models that can reason across experimental data, simulations, and text—but we’re still far from that level of reliability or interpretability.
  • AI for experimental control: AI-driven experimental control is enabling autonomous, adaptive laboratories where algorithms plan and adjust experiments in real time, as shown by recent self-driving lab demonstrations (from AI-guided organic synthesis to autonomous materials screening in both academic and industrial settings) that optimized reaction conditions and achieved successful syntheses with minimal human intervention. In the near term, integrating AI controllers with robotics and advanced sensors will allow experiments such as multi-step chemical syntheses or high-throughput materials discovery to self-optimize based on interim results, dramatically accelerating research cycles. Ultimately, fully autonomous laboratories could execute complex research programs—formulating hypotheses, running iterative experiments, and interpreting results—vastly speeding discovery in fields from drug development to energy storage, beyond the pace of human-guided work. It remains to be seen how much the role of human-in-the-loop can be minimized or even eliminated, and if that is something worthy of pursuing due to possible hazards with synthesizing chemicals that could be dangerous.
@@
  • Develop models for new areas, such as new large language models (LLMs) for chemistry and biology that can design proteins, catalysts, small molecules, large molecules, and materials with desired functions that can be tuned for different environments, revolutionizing applications in medicine, bioenergy, sustainability, and chemical synthesis. These models should go beyond static structure paradigm to embed thermodynamic and kinetic considerations directly into the generation process. Doing so could allow for more targeted design in out-of-equilibrium and complex environments.
@@
  • Accelerate the rate of discovery by incorporating methods to ensure the synthesizability of candidates into discovery workflows and/or identify synthesizable analogs.
    Models should not just predict exciting molecules but be able to assess chemical novelty, while also suggesting how to make them under realistic lab conditions. Incorporating synthesis constraints early could prevent costly dead ends downstream.
@@ 3.3.1 AI in the Context of DMR
In general, DMR research, like CHE (Sec. 3.2) operates in the small-data regime. The reason for this is that materials research is exceedingly heterogeneous, with research spanning from perovskites to transporter proteins, and the interesting materials are—almost by definition—rare, meaning that there is sparse data in their vicinity. Furthermore, the details of composition, microstructure, and (in many cases) processing history are all important for understanding and predicting materials properties. Soft materials must also contend with entropic considerations, wherein the properties of the material are not dictated by a single structure but an ensemble of configurations. Furthermore, interesting materials with desirable properties may not be at equilibrium, meaning that it is often important to account for kinetic considerations and dynamical processing in AI/ML models. This can also mean incorporating factors from actual growth, synthesis, and fabrication processes—such as temperature, pressure, and fabrication resolution constraints—to better reflect real-world scenarios.
@@ 3.4.2 How AI is Advancing DMS
  • AI in scientific computing: Complementing theory and experiment, scientific computing has become the third pillar of science. Integrating AI has led to the development of the vibrant field of scientific machine learning (SciML) that supports emerging technologies such as digital twins by leveraging techniques from AI, applied mathematics, computational mathematics, statistics, and probability. Some key themes in this area are:
    • Discover, refine, and optimize numerical algorithms: AI helps with searching for novel numerical schemes, optimizing solver configurations, and learning adaptive discretizations and iterative methods. These efforts suggest a future in which AI is not merely a tool for modeling but aids algorithmic design, informed by mathematical insights from approximation theory, optimization, and numerical analysis. For example, Defense Advanced Research Projects Agency (DARPA)'s ``Mathematics for the Discovery of Algorithms and Architectures'' (DIAL) program [^30] https://sam.gov/workspace/contract/opp/8e9428a0783e483f954d93555dd8cc77/view is developing methods to discover new generalized numerical algorithms in a principled way. Similarly, their ``The Right Space'' (TRS) program [^31] https://sam.gov/opp/104a7ef9ccef4fa1b47aa9eaf4c89d0a/view is seeking to create and field new tools for discovering useful mathematical transformations and establish the domain of ``algorithms for numerics for singularities.''
@@
  • AI-assisted theorem proving: Recent advances toward using AI tools for formal reasoning and as ``theorem prover copilots'' helps mathematicians formalize proofs and enables interactive proof verification, such as in co-pilots for Lean [^32] https://lean-lang.org/ and advances in SAT solvers. The use of proof assistants can reveal deep mathematical connections and has, for example, led to new logical insight about the notion of equality and unexpected connections between equality and the field of homotopy theory. Homotopy type theory is considered a relevant complement to existing axiomatic foundations of proof assistants, especially for proofs involving diagrams or higher dimensional structures.
  • AI in mathematical biology: AI has revolutionized protein structure prediction and offers powerful tools for modeling complex biological systems. While traditional computational methods such as homology modeling, protein threading, and ab initio modeling have long been used to predict protein structures from amino acid sequences, they faced challenges in accuracy, scalability, and generalization to novel proteins. AI, especially deep learning, enable efficient search of high-dimensional parameter spaces and extract patterns from massive sequence databases, which dramatically improves predictive power and robustness. State-of-the-art models such as AlphaFold2, [^33] https://github.com/google-deepmind/alphafold RoseTTAFold, [^34] https://github.com/RosettaCommons/RoseTTAFold and ProteinBERT [^35] https://github.com/nadavbra/protein_bert have set new standards for protein structure prediction. Notably, AlphaFold2 achieved near-experimental accuracy in the CASP14 competition and demonstrated utility across the life sciences. These models enable a deeper understanding of protein function, molecular interactions, and therapeutic target identification by reliably inferring 3D structures, and are critical to interpreting the vast amounts of sequence data generated by large-scale genomics and proteomics initiatives.
@@ 3.4.3 How DMS is Advancing AI
  • Training dynamics and optimization: Almost all learning tasks are formulated as optimization problems. Consequently, modern mathematical tools resulting from DMS research have become essential for understanding and enhancing the training dynamics of DNNs. First-order optimization methods, including stochastic gradient descent, momentum-based variants, and adaptive algorithms such as Adam, build upon classical optimization research. [^36] https://doi.org/10.1137/16M1080173
@@ 3.4.4 Opportunities in AI+DMS
    • Topological data analysis (TDA): Built upon advanced mathematical concepts from algebraic topology, TDA reveals robust geometric and structural insights in complex, high-dimensional data. Using tools such as persistent homology, simplicial complexes, and topological invariants, TDA provides AI systems with meaningful shape-based features that remain stable under noise and deformation. These methods complement statistical and optimization-based approaches and have proven particularly effective in improving the interpretability and robustness of AI models. Recent advances further integrate TDA directly into neural network architectures through differentiable persistence modules, guiding model training toward capturing intrinsic topological features and leading to more reliable, interpretable AI solutions.
@@
    • Digital twins: Mathematical abstraction is also fundamental in the design of digital twins. DMS research on inverse problems, model reduction, UQ, structural identifiability, and optimal control enriches AI tools and supports the creation and refinement of digital twins.
@@ 3.5.2 How AI is Advancing PHY
The broad spectrum of approaches and needs within the physics community have already led to innovative research in AI+Physics upon which to build. AI has been used to advance the analysis of large datasets, enhance pattern recognition, improve the detection and removal of noise artifacts in data, and revolutionize simulations. Below are some specific examples from each of the physics subdomains that demonstrate success interfacing with AI. We emphasize that this is not an exhaustive list of examples, but rather case studies drawn from the workshop participants.
@@
  • Elementary particle physics has been a voracious user of AI for collider experiments. Existing modules in the collider analysis pipeline have been replaced with AI-powered alternatives, which have been delivering broad performance gains. Deep-learning-based AI has enabled high-dimensional statistical analyses, allowing physicists to leverage far more information than was previously possible, without relying solely on summary statistics. More recently, anomaly detection methods have been deployed to support the discovery of physics beyond the Standard Model. Anomaly detection represents a fundamental change in the way data is collected, processed, and interpreted, which would not be possible without AI. These efforts have fostered interdisciplinary collaboration among theorists, experimentalists, ML researchers, and hardware specialists. One notable outcome is the development of real-time AI systems for large-scale experiments, where AI can now be deployed at the LHC to process data with just 50 nanosecond latency. Additional uses of AI in high energy physics include event reconstruction, object identification, SBI, neural importance sampling, surrogate modeling, full phase-space unfolding, and more. [^37] https://iml-wg.github.io/HEPML-LivingReview/
@@ 3.5.3 How PHY is Advancing AI
As mentioned previously, a notable example of physics advancing AI can be seen in the 2024 Nobel Prize in Physics, which went to Hopfield & Hinton for the development of deep learning by combining statistical physics with computational neuroscience. [^38] https://www.nobelprize.org/prizes/physics/2024/summary/ Building on that, the development of diffusion models was inspired by concepts from statistical physics. Here are additional examples where progress has been made in physics influencing AI:
@@ 3.5.4 Opportunities in AI+PHY
  • Foundation models: The development of Large Physics Models—foundation models trained across diverse physics domains—represents an opportunity to accelerate data analysis and scientific discovery in experiments such as the LHC, LIGO, EIC, and beyond. By capturing shared structure across datasets, these models could enable transfer learning, improve sample efficiency, and provide unified frameworks for tasks ranging from simulation and reconstruction to anomaly detection and inference.
@@ Acknowledgments
Acronyms
ADS
Astrophysics Data System
AGI
artificial general intelligence
AI
artificial intelligence
API
application programming interface
AST
(NSF Division of) Astronomical Sciences
BDT
boosted decision tree
CCAS
Center for Computer-Assisted Synthesis
CDS
Strasbourg astronomical Data Center
CHE
(NSF Division of) Chemistry
CIF
Crystallographic Information Framework
CISE
(NSF Directorate for) Computer and Information Science and Engineering
CNN
convolutional neural network
CS
computer science
CSSI
Cyberinfrastructure for Sustained Scientific Innovation
DARPA
Defense Advanced Research Projects Agency
DFT
density functional theory
DLM
data lifecycle management
DMR
(NSF Division of) Materials Research
DMS
(NSF Division of) Mathematical Sciences
DNN
deep neural network
DOE
Department of Energy
DUNE
Deep Underground Neutrino Experiment
EIC
Electron-Ion Collider
GAN
generative adversarial network
GNN
graphic neural network
GPU
graphics processing unit
GRMHD
general-relativistic magnetohydrodynamic
HL-LHC
High-luminosity LHC
HNN
Hamiltonian neural network
HPC
high-performance computing
IAIFI
Institute for Artificial Intelligence and Fundamental Interactions
ICSD
Inorganic Crystal Structure Database
KG
knowledge graph
LHC
Large Hadron Collider
LIGO
Laser Interferometer Gravitational-wave Observatory
LISA
Laser Interferometer Space Antenna
LLM
large language model
LNN
Lagrangian neural network
LoRA
Low Rank Adaptation
MaaS
Model as a Service
MFAI
Mathematical Foundations of AI
ML
machine learning
MLIP
machine-learned interatomic potential
MoDL
Mathematical and Scientific Foundations of Deep Learning
MPS
(NSF Directorate for) Mathematical and Physical Sciences
NAIRR
National AI Research Resource
NASA
National Aeronautics and Space Administration
NeurIPS
Conference on Neural Information Processing Systems
NF
normalizing flow
NISQ
noisy intermediate-scale quantum
NITMB
National Institute for Theory and Mathematics in Biology
NN
neural network
NSF
National Science Foundation
NTK
neural tangent kernel
ODE
ordinary differential equation
PDB
Protein Data Bank
PDE
partial differential equation
PHY
(NSF Division of) Physics
PI
principal investigators
PINN
physics-informed neural network
PNT
positioning, navigation, and timing
QFT
quantum field theory
RAG
retrieval augmented generation
RAGA
retrieval augmented generation assessment
RKHS
reproducing kernel Hilbert space
RL
reinforcement learning
RMT
random matrix theory
SAT
Boolean satisfiability problem
SBI
simulation-based inference
SciML
scientific machine learning
SDE
stochastic differential equation
SMT
satisfiability modulo theories
STEM
science, technology, engineering, and mathematics
TDA
topological data analysis
UQ
uncertainty quantification
VAE
variational autoencoder

### v3

@@ 3.2.1 AI in the Context of CHE
Chemistry and AI have long been working together, creating an exchange that is reshaping both fields. On the one hand, AI is helping chemists predict reactions, design molecules, and analyze complex datasets with speed and precision—accelerating discovery across chemical synthesis, drug development, materials science, and catalysis. On the other hand, chemistry is pushing the frontiers of AI. Theoretical chemistry has inspired some of today’s most powerful generative models, including diffusion models which were directly influenced by non-equilibrium thermodynamics. In fact, chemistry has provided some of AI’s most celebrated successes; notably, AlphaFold earned the 2024 Nobel Prize in Chemistry [undefb] for its breakthrough in protein structure prediction and protein design. These examples highlight a growing, bidirectional relationship: AI is transforming how we do chemistry, and chemistry continues to shape and challenge the future of AI itself.
@@ 3.2.2 How AI is Advancing CHE
  • Computational chemistry increasingly uses AI tools to predict molecular behavior and accelerate simulations, particularly for complex systems such as protein dynamics or excited-state reactions in a complex medium [undefah]. These often involve developing surrogate models to replace expensive simulation methods based on quantum mechanics. Beyond speed-ups, AI is also aiding in automated design of coarse-grained models, identifying low-dimensional collective variables that capture essential system dynamics, and enhancing sampling of rare events that are otherwise inaccessible to conventional simulations. These different sub-disciplines of computational chemistry are now rapidly being transformed with the thoughtful use of AI, often deeply integrated with statistical mechanics.
@@
  • Interatomic potentials and atomistic simulation studies have been transformed by machine learning, enabling first-principles accuracy with far greater computational efficiency. While traditional first-principles approaches like density functional theory (DFT) are limited to small systems (hundreds of atoms) and short timescales (picoseconds), machine-learned interatomic potentials (MLIPs) make it possible to simulate tens of thousands of atoms over nanoseconds or longer [undefai]. These advances provide efficient and accurate ways to predict reaction pathways and transition states, rapidly screen molecular libraries, and design novel hypothetical materials with targeted properties for experimental synthesis.
@@ 3.2.4 Opportunities in AI+CHE
  • Foundation models: Foundation models have shown promise in handling a range of chemistry tasks, from property prediction to retrosynthesis and even literature mining. Fine-tuning them for synthesis, materials, or protein design could streamline workflows, but questions remain about how well they generalize to out-of-distribution chemistry or rare reactions. The longer-term opportunity is to build models that can reason across experimental data, simulations, and text—but we are still far from that level of reliability or interpretability.
  • AI for experimental control: AI-driven experimental control is enabling autonomous, adaptive laboratories where algorithms plan and adjust experiments in real time, as shown by recent self-driving lab demonstrations (from AI-guided organic synthesis to autonomous materials screening in both academic and industrial settings) that optimized reaction conditions and achieved successful syntheses with minimal human intervention. In the near term, integrating AI controllers with robotics and advanced sensors will allow experiments such as multi-step chemical syntheses or high-throughput materials discovery to self-optimize based on interim results, dramatically accelerating research cycles. Ultimately, fully autonomous laboratories could execute complex research programs—formulating hypotheses, running iterative experiments, and interpreting results—vastly speeding discovery in fields from drug development to energy storage, beyond the pace of human-guided work. It remains to be seen how much the role of human-in-the-loop can be minimized or even eliminated, and if that is something worthy of pursuing due to possible hazards with synthesizing chemicals that could be dangerous. In particular, it will be important to bear in mind the requirement for nucleic acid synthesis screening and ``safety-by-design'' protocols in federally-funded automated laboratories.
@@
  • Develop models for new areas, such as new large language models (LLMs) for chemistry and biology that can design proteins, catalysts, small molecules, large molecules, and materials with desired functions that can be tuned for different environments, revolutionizing applications in medicine, bioenergy, sustainability, and chemical synthesis [undefaj]. These models should go beyond static structure paradigm to embed thermodynamic and kinetic considerations directly into the generation process. Doing so could allow for more targeted design in out-of-equilibrium and complex environments.
@@
  • Accelerate the rate of discovery by incorporating methods to ensure the synthesizability of candidates into discovery workflows and/or identify synthesizable analogs. Models should not just predict exciting molecules but be able to assess chemical novelty, while also suggesting how to make them under realistic lab conditions. Incorporating synthesis constraints early could prevent costly dead ends downstream.
@@ 3.3.1 AI in the Context of DMR
In general, DMR research, like CHE (Sec. 3.2), operates in the small-data regime. The reason for this is that materials research is exceedingly heterogeneous, with research spanning from perovskites to transporter proteins, and the interesting materials are—almost by definition—rare, meaning that there is sparse data in their vicinity. Furthermore, the details of composition, microstructure, and (in many cases) processing history are all important for understanding and predicting materials properties. Soft materials must also contend with entropic considerations, wherein the properties of the material are not dictated by a single structure but an ensemble of configurations. Furthermore, interesting materials with desirable properties may not be at equilibrium, meaning that it is often important to account for kinetic considerations and dynamical processing in AI/ML models. This can also mean incorporating factors from actual growth, synthesis, and fabrication processes—such as temperature, pressure, and fabrication resolution constraints—to better reflect real-world scenarios.
@@ 3.4.2 How AI is Advancing DMS
  • AI in scientific computing: Complementing theory and experiment, scientific computing has become the third pillar of science. Integrating AI has led to the development of the vibrant field of scientific machine learning (SciML) that supports emerging technologies such as digital twins by leveraging techniques from AI, applied mathematics, computational mathematics, statistics, and probability [undeff]. Some key themes in this area are:
    • Discover, refine, and optimize numerical algorithms: AI helps with searching for novel numerical schemes, optimizing solver configurations, and learning adaptive discretizations and iterative methods. These efforts suggest a future in which AI is not merely a tool for modeling but aids algorithmic design, informed by mathematical insights from approximation theory, optimization, and numerical analysis. For example, Defense Advanced Research Projects Agency (DARPA)'s ``Mathematics for the Discovery of Algorithms and Architectures'' (DIAL) program [undefak] is developing methods to discover new generalized numerical algorithms in a principled way. Similarly, their ``The Right Space'' (TRS) program [undefal] is seeking to create and field new tools for discovering useful mathematical transformations and establish the domain of ``algorithms for numerics for singularities.''
@@
  • AI-assisted theorem proving: Recent advances toward using AI tools for formal reasoning and as ``theorem prover copilots'' helps mathematicians formalize proofs and enables interactive proof verification, such as in co-pilots for Lean and advances in SAT solvers [undefam]. The use of proof assistants can reveal deep mathematical connections and has, for example, led to new logical insight about the notion of equality and unexpected connections between equality and the field of homotopy theory. Homotopy type theory is considered a relevant complement to existing axiomatic foundations of proof assistants, especially for proofs involving diagrams or higher dimensional structures.
  • AI in mathematical biology: AI has revolutionized protein structure prediction and offers powerful tools for modeling complex biological systems. While traditional computational methods such as homology modeling, protein threading, and ab initio modeling have long been used to predict protein structures from amino acid sequences, they faced challenges in accuracy, scalability, and generalization to novel proteins. AI, especially deep learning, enable efficient search of high-dimensional parameter spaces and extract patterns from massive sequence databases, which dramatically improves predictive power and robustness. State-of-the-art models such as AlphaFold2 [undefan], RoseTTAFold [undefao], and ProteinBERT [undefap] have set new standards for protein structure prediction. Notably, AlphaFold2 achieved near-experimental accuracy in the CASP14 competition and demonstrated utility across the life sciences. These models enable a deeper understanding of protein function, molecular interactions, and therapeutic target identification by reliably inferring 3D structures, and are critical to interpreting the vast amounts of sequence data generated by large-scale genomics and proteomics initiatives.
@@ 3.4.3 How DMS is Advancing AI
  • Training dynamics and optimization: Almost all learning tasks are formulated as optimization problems. Consequently, modern mathematical tools resulting from DMS research have become essential for understanding and enhancing the training dynamics of DNNs. First-order optimization methods, including stochastic gradient descent, momentum-based variants, and adaptive algorithms such as Adam, build upon classical optimization research [undefaq].
@@ 3.4.4 Opportunities in AI+DMS
    • Topological data analysis (TDA): Built upon advanced mathematical concepts from algebraic topology, TDA reveals robust geometric and structural insights in complex, high-dimensional data [undefar]. Using tools such as persistent homology, simplicial complexes, and topological invariants, TDA provides AI systems with meaningful shape-based features that remain stable under noise and deformation. These methods complement statistical and optimization-based approaches and have proven particularly effective in improving the interpretability and robustness of AI models. Recent advances further integrate TDA directly into neural network architectures through differentiable persistence modules, guiding model training toward capturing intrinsic topological features and leading to more reliable, interpretable AI solutions.
@@
    • Digital twins: Mathematical abstraction is also fundamental in the design of digital twins [undefas]. DMS research on inverse problems, model reduction, UQ, structural identifiability, and optimal control enriches AI tools and supports the creation and refinement of digital twins.
@@ 3.5.2 How AI is Advancing PHY
The broad spectrum of approaches and needs within the physics community have already led to innovative research in AI+Physics upon which to build [undefat]. AI has been used to advance the analysis of large datasets, enhance pattern recognition, improve the detection and removal of noise artifacts in data, and revolutionize simulations. Below are some specific examples from each of the physics subdomains that demonstrate success interfacing with AI. We emphasize that this is not an exhaustive list of examples, but rather case studies drawn from the workshop participants.
@@
  • Elementary particle physics has been a voracious user of AI for collider experiments. Existing modules in the collider analysis pipeline have been replaced with AI-powered alternatives, which have been delivering broad performance gains. Deep-learning-based AI has enabled high-dimensional statistical analyses, allowing physicists to leverage far more information than was previously possible, without relying solely on summary statistics. More recently, anomaly detection methods have been deployed to support the discovery of physics beyond the Standard Model. Anomaly detection represents a fundamental change in the way data is collected, processed, and interpreted, which would not be possible without AI. These efforts have fostered interdisciplinary collaboration among theorists, experimentalists, ML researchers, and hardware specialists. One notable outcome is the development of real-time AI systems for large-scale experiments, where AI can now be deployed at the LHC to process data with just 50 nanosecond latency. Additional uses of AI in high energy physics include event reconstruction, object identification, SBI, neural importance sampling, surrogate modeling, full phase-space unfolding, and more [undefau].
@@ 3.5.3 How PHY is Advancing AI
As mentioned previously, a notable example of physics advancing AI can be seen in the 2024 Nobel Prize in Physics, which went to Hopfield & Hinton for the development of deep learning by combining statistical physics with computational neuroscience [undefav]. Building on that, the development of diffusion models was inspired by concepts from statistical physics. Here are additional examples where progress has been made in physics influencing AI:
@@ 3.5.4 Opportunities in AI+PHY
  • Foundation models: The development of Large Physics Models—foundation models trained across diverse physics domains—represents an opportunity to accelerate data analysis and scientific discovery in experiments such as the LHC, LIGO, EIC, and beyond. This could be a single monolithic model (with an interface similar to a commercial general-purpose LLM), or a system of interoperable domain-specific models (offering a higher degree of customization for varied scientific workflows). By capturing shared structure across datasets, these models could enable transfer learning, improve sample efficiency, and provide unified frameworks for tasks ranging from simulation and reconstruction to anomaly detection and inference.